# Learning 3-Manifold Triangulations: Pachner Graph Network Analysis & Machine Learning IsoSigs

### v1

@@ 4.2 Differentiating Knots
Table 4: The knot IsoSigs used to start the generation of Pachner graphs using only $2-3$ and $3-2$ moves. Each knot was used to construct a knot-complement 3-manifold, which were triangulated to give a 1-vertex triangulation. The respective IsoSigs are listed in this table, along with the number of tetrahedra in this initial triangulation.
@@
The results show comparable performances between knots, however with worse maximum performances compared to the general 3-manifold learning in §4.1. Some pairings had close to no learning, better exemplified by the near-zero MCC scores for the (FigureEight,$5_{2}$) and the ($5_{2}$,DT $6a_{3}$) pairings. Generally, as shown by the images, the colours are darker closer to the diagonal, indicating worse performance, and suggesting that in general the more disparate the crossing structure of the knot the easier the NNs can distinguish them. These accuracies obtained in the task of comparing with the unknot are by far less good than those shown for this specific task in [37].
@@ 4.2.1 Supplementing with Dehn Surgery
Focus is put on the differentiation between the Unknot and Trefoil binary classification problem, which for the knot complement $3$-manifolds classified with accuracy 0.632 and MCC 0.265, neither particularly impressive scores. In this extension, the initial knot complement triangulation IsoSigs, as given in Table 4, had either 0-surgery or 1-surgery performed on them, to produce further $3$-manifolds whose IsoSigs are given in Table 5. We remind the reader that the 00 and $1$ surgeries on the unknot are respectively $S^{2}\times S^{1}$ and $S^{3}$ and on the left handed trefoil they are two Seifert fibered manifolds, respectively $(S^{2},(2,1),(3,1),(6,-5))$ and $(S^{2},(2,1),(3,1),(7,-6))$ (the latter not being the Poincaré homology sphere, which is the $1$ surgery over the right handed trefoil). Each of these was used to seed Pachner graph generation to generate databases of 2000 length 30 IsoSigs. Noting that in this subsection length 30 was used again instead of the length 25 as used in §4.2, this is because the initial IsoSigs after surgery lead to length distributions with very few length 25 IsoSigs [^12] This is also why just these two simplest knots are considered since the length distributions for some knot complements after surgery did not even include any length 25 or 30 IsoSigs.. Another comment worth emphasising is that whilst the IsoSigs in Table 5 have a range of initial number of tetrahedra, since only a fixed length IsoSig is taken from the graph for ML these IsoSigs will also correspond to triangulations with the same number of tetrahedra. Additionally, it is worth noting that the initial IsoSig from 0-surgery of the Unknot is a triangulation with 3-vertices (whilst the others are still all 1-vertex), this means this Pachner graph is a no longer a 1-vertex Pachner graph.
@@ Appendix B Pachner Graphs
$S^{2}\times S^{1}$ Depth 3
$S^{2}\times S^{1}$ Depth 4
$\mathbb{RP}^{3}$ Depth 3
$\mathbb{RP}^{3}$ Depth 4
$L(7,1)$ Depth 3
$L(7,1)$ Depth 4
$L(7,2)$ Depth 3
$L(7,2)$ Depth 4
$T^{3}$ Depth 3
$T^{3}$ Depth 4
PHS Depth 3
PHS Depth 4
$H_{SC}$ Depth 1
$H_{SC}$ Depth 2
$S^{3}$ depth 5

### v2

@@ 4.2 Differentiating Knots
Table 4: The knot IsoSigs used to start the generation of Pachner graphs using only $2-3$ and $3-2$ moves. Each knot was used to construct a knot-complement 3-manifold, which was then triangulated to give a 1-vertex triangulation. The respective IsoSigs are listed in this table, along with the number of tetrahedra in this initial triangulation.
@@
The results show comparable performances between knots, however with worse maximum performances compared to the general 3-manifold learning in §4.1. Some pairings had close to no learning, better exemplified by the near-zero MCC scores for the (FigureEight, $5_{2}$) and the ($5_{2}$, DT $6a_{3}$) pairings. Generally, as shown by the images, the colours are darker closer to the diagonal, indicating worse performance, and suggesting that in general the more disparate the crossing structure of the knot the easier the NNs can distinguish them. Accuracies for identifying the Unknot are not as high as in the focused Unknot study in [44].
@@ 4.2.1 Supplementing with Dehn Surgery
Focus is put on the differentiation between the Unknot and Trefoil binary classification problem, which for the knot complement $3$-manifolds classified with accuracy 0.632 and MCC 0.265, neither particularly impressive scores. In this extension, the initial knot complement triangulation IsoSigs, as given in Table 4, had either 0-surgery or 1-surgery performed on them, to produce further $3$-manifolds whose IsoSigs are given in Table 5. We remind the reader that the 00 and $1$ surgeries on the Unknot are respectively $S^{2}\times S^{1}$ and $S^{3}$ and on the left-handed Trefoil they are two Seifert fibered manifolds, respectively $(S^{2},(2,1),(3,1),(6,-5))$ and $(S^{2},(2,1),(3,1),(7,-6))$ (the latter not being the Poincaré homology sphere, which is the $1$ surgery over the right-handed Trefoil). Each of these was used to seed Pachner graph generation to generate databases of 2000 length 30 IsoSigs. Noting that in this subsection length 30 was used again instead of the length 25 as used in §4.2, this is because the initial IsoSigs after surgery lead to length distributions with very few length 25 IsoSigs [^12] This is also why just these two simplest knots are considered since the length distributions for some knot complements after surgery did not even include any length 25 or 30 IsoSigs.. Another comment worth emphasising is that whilst the IsoSigs in Table 5 have a range of initial number of tetrahedra, since only a fixed length IsoSig is taken from the graph for ML these IsoSigs will also correspond to triangulations with the same number of tetrahedra. Additionally, it is worth noting that the initial IsoSig from 0-surgery of the Unknot is a triangulation with 3-vertices (whilst the others are still all 1-vertex), this means this Pachner graph is a no longer a 1-vertex Pachner graph.
@@ Appendix B Pachner Graphs
Figure 17: $S^{2}\times S^{1}$ Depth 3
Figure 18: $S^{2}\times S^{1}$ Depth 4
Figure 19: $\mathbb{RP}^{3}$ Depth 3
Figure 20: $\mathbb{RP}^{3}$ Depth 4
Figure 21: $L(7,1)$ Depth 3
Figure 22: $L(7,1)$ Depth 4
Figure 23: $L(7,2)$ Depth 3
Figure 24: $L(7,2)$ Depth 4
Figure 25: $T^{3}$ Depth 3
Figure 26: $T^{3}$ Depth 4
Figure 27: PHS Depth 3
Figure 28: PHS Depth 4
Figure 29: $H_{SC}$ Depth 1
Figure 30: $H_{SC}$ Depth 2
Figure 31: $S^{3}$ depth 5